OpenStudy (anonymous):

The demand for plastic brownie dishes is given below where q represents the number of brownie dishes that can be sold at a price of p q(p)=363609-(p+1)^2 1. Use q(p) to determine the lowest price at which it will be unable to sell any dishes (basically what formula do I need to use to solve this?)

OpenStudy (anonymous):

sorry idk

OpenStudy (anonymous):

Set the equation equal to 0, because q(p) is how many dishes can be sold at a certain price, and you want to find the price where you can sell 0 dishes (no dishes)

OpenStudy (anonymous):

okay leme try

OpenStudy (anonymous):

Set the equation equal to 0, because q(p) is how many dishes can be sold at a certain price, and you want to find the price where you can sell 0 dishes (no dishes)

OpenStudy (anonymous):

would it start out like 363609=(p+1)^2? or should i simplify out the (p+1)^2 to p^2+2p+1?

OpenStudy (anonymous):

Yes, expand it into a quadratic equation and solve it as you would

OpenStudy (anonymous):

so it would be 363609=p^2+2p+1?

OpenStudy (anonymous):

and then solve for p?

OpenStudy (anonymous):

Move all the terms over to one side first and set it equal to 0

OpenStudy (anonymous):

-p^2+2p=363608?

OpenStudy (anonymous):

No, 0 = p^2 + 2p - 363607

OpenStudy (anonymous):

would that be the answer then?

OpenStudy (anonymous):

Wait, let me check the original equation first

OpenStudy (anonymous):

\[ 0 = 363,609 - (p+1)^2 = 363,609 - p^2 + 2p + 1 = p^2 - 2p - 363,610\]

OpenStudy (anonymous):

so 0=p^2-2p-363610?

OpenStudy (anonymous):

how would i solve for p then sense it is in two separate forms?

OpenStudy (anonymous):

You can use the quadratic equation, and the price should be a positive number

OpenStudy (anonymous):

[2\pm \sqrt{1454444}/2\]?

OpenStudy (anonymous):

\[2\pm \sqrt{1454444}/2\]?

OpenStudy (anonymous):

\[\frac{-(-2) \pm \sqrt{(-2)^2 - 4(1)(-363,610)}}{2(1)} = \frac{2 \pm \sqrt{4 - (-1,454,440)}}{2} = \\ \frac{2\pm \sqrt{1454444}}{2} = 1\pm \sqrt{363,611}\] And 1 - sqrt(363,611) is going to be negative so you choose the 1+ sqrt(363,611) value as 'p'

OpenStudy (anonymous):

so your saying the answer would be 604 then?

OpenStudy (anonymous):

Rounded, yes

OpenStudy (anonymous):

its saying that isnt correct ;/

OpenStudy (anonymous):

Let me look back

OpenStudy (anonymous):

Does the question say \[q(p) = 363,609 - (p + 1)^2\] or \[q(p) = 363,609 \cdot -(p + 1)^2\]

OpenStudy (anonymous):

the first one u put

OpenStudy (anonymous):

In any case, for the first one, I think I made an error in sign distribution while I was typing it...I worked it out on paper and it came out differently; the solutions are -604 and 602 so it's 602 to satisfy the original equation

OpenStudy (anonymous):

correct thank you so much!